If a black body is heated from 27°C to 927°C. then the ratio of radiation emitted will be
(a) 1: 4
(b) 1: 16
(c) 1: 8
(d) 1: 256

Grade:Upto college level

1 Answers

Devansh Jain
6 years ago
(d) 1:256 
27+273= 300K
927+273=1200K
we now have, 
                     ratio of temperatures= 300/1200
                                                       = ¼
                      therefore, ratio of radiation emitted = 1^4/4^4        [(T1)^4/(T2)^4]
                                                                               = 1/256